OpenStudy (anonymous):

for the equation -4y=8x what is the constant of variation

OpenStudy (anonymous):

It would be -2 because the equation for the constant variation is y=kx k being the constant variation so you just put the equation in constant cariation form -4y=8x -4/(-4y)= -4/(8x) y=-2x y=kx which means that k (constant variation) is -2 :D

OpenStudy (anonymous):

oh ok thanks :)

OpenStudy (anonymous):

no problem :D

OpenStudy (anonymous):

did you understand it ?

OpenStudy (anonymous):

yes i did :)